OpenStudy (anonymous):

f' and f'' of : f(x)= (x^3)/3 - 1/x ??

OpenStudy (lncognlto):

Okay, so first we re-write it: f(x)= (x^3)/3 - 1/x therefore f(x) = (x^3)/3 - x^-1 tf, following the rules of differentiation, f'(x) = (3x^2)/3 - (-1)x^-2... tf f'(x) = x^2 + x^-2 tf f'(x) = x^2 + 1/x^2

OpenStudy (lncognlto):

Now, copying what I did there, what would f''(x) be...?

OpenStudy (anonymous):

yes I did the first part, in fact, I have difficulties with the other part...

OpenStudy (anonymous):

(f'')

OpenStudy (anonymous):

That'S what I did but it's not the goood answer

OpenStudy (lncognlto):

Okay, I see what happened. You are correct, to a point. xD f"(x) does indeed equal 2x - 2x^-3 but then the next step is that f"(x) = 2x-2/x^3, not (2x-2)/x^3.

OpenStudy (anonymous):

whats the difference?

OpenStudy (anonymous):

the real answer is :( 2(x^2 + 1)(x-1)(x+1) ) / x^3

OpenStudy (lncognlto):

O.o

OpenStudy (lncognlto):

I'm flummoxed. LOL

OpenStudy (lncognlto):

But let me think...

OpenStudy (lncognlto):

OHHHH

OpenStudy (lncognlto):

I have no idea why they wrote it like this, but if we expand 2(x^2 + 1)(x-1)(x+1) ) / x^3, we eventually come out at 2x-2/x^3. LOL

OpenStudy (anonymous):

okay haha thanls :P

OpenStudy (anonymous):